# Improving the fact-checking performance of language models by relying on their entailment ability

### v1

@@ Abstract
Automated fact-checking is a crucial task in this digital age. To verify a claim, current approaches majorly follow one of two strategies i.e. (i) relying on embedded knowledge of language models, and (ii) fine-tuning them with evidence pieces. While the former can make systems to hallucinate, the later have not been very successful till date. The primary reason behind this is that fact verification is a complex process. Language models have to parse through multiple pieces of evidence before making a prediction. Further, the evidence pieces often contradict each other. This makes the reasoning process even more complex. We proposed a simple yet effective approach where we relied on entailment and the generative ability of language models to produce “supporting” and “refuting” justifications (for the truthfulness of a claim). We trained language models based on these justifications and achieved superior results. Apart from that, we did a systematic comparison of different prompting and fine-tuning strategies, as it is currently lacking in the literature. Some of our observations are: (i) training language models with raw evidence sentences registered an improvement up to 8.20% in macro-F1, over the best performing baseline for the RAW-FC dataset, (ii) similarly, training language models with prompted claim-evidence understanding (TBE-2) registered an improvement (with a margin up to 16.39%) over the baselines for the same dataset, (iii) training language models with entailed justifications (TBE-3) outperformed the baselines by a huge margin (up to 28.57% and 44.26% for LIAR-RAW and RAW-FC, respectively). We have shared our code repository to reproduce the results.
@@ 1 Introduction:
The spread of misinformation on the internet has grown to be a pressing social issue. Its consequences have manifested across social, political and commercial domains Mozur (2018); Fisher et al. (2016); Allcott and Gentzkow (2017); Burki (2019); Aghababaeian et al. (2020). To counter the spread, institutional interventions doing manual fact-checking have gained momentum. For example, the International Fact-Checking Network (IFCN), started in 2015, works with over 170 fact-checking organisations and websites worldwide (as of July 2024  [^1] https://www.poynter.org/ifcn/). However, manually verifying facts and detecting misinformation is a slow and costly process. Human experts struggle to keep up with the pace of rapid spread. To overcome this, the research community have been trying to automate the process. Their interest can be gauged by the fact that more than twelve hundred research articles and more than fifty datasets were published on this topic Alnabhan and Branco (2024); Guo et al. (2022). A detailed list of such related works is reported in the Appendix A.
@@
Doing automated fact-checking at scale is still an open challenge. Most of the proposed approaches relied on language models Shu et al. (2022); Yang et al. (2022a); Yue et al. (2023); Choi and Ferrara (2024a). Researchers have either (i) relied on their embedded knowledge or (ii) used evidence collected from various sources to predict the veracity. For example, Pan et al. (2021); Zeng and Gao (2023) reported the performance of zero-shot or few-shot prompting for this task. Dhuliawala et al. (2024) showcased that integrating reasoning steps through chain-of-thought prompting reduced model hallucinations. Others relied on external tools such as web search Galitsky (2025); Zhang and Gao (2023) to retrieve the evidence and do fact verification. However, their accuracy is far behind to deploy them for practical purposes. The primary reason behind this is that fact verification is difficult because language models need to look at multiple pieces of information to decide what is true. These pieces often disagree with each other, which makes the reasoning process even harder. To counter this, Yue et al. (2024) segregated the retrieved evidence into ‘supporting’ and ‘opposing’ arguments, allowing the model to consider each perspective independently, and then they verified each claim using few-shot prompting. On a philosophically similar line, Wang et al. (2024a) introduced a module to extract top-k evidence sets stating the claim to be ‘true’ and ‘false’ based on attention scores. They produced ‘true’ and ‘false’ justifications by promoting language models based on evidence sentences, and further trained the language models to produce veracity by training LLMs. We, on the other hand, proposed a simple yet effective approach, where we relied on (i) the entailment ability of language models to classify if evidence sentences are supporting or refuting a given claim and (ii) the generative ability of language models to produce supporting and refuting justifications. A schematic diagram of our proposed approach is presented in Figure 17 (in the Appendix). We believe the simplicity of our approach will allow us to easily deploy our systems in the real-world. Apart from that, we also found that a systematic comparison of different prompting and fine-tuning strategies is lacking in the literature. To fill this research gap, we designed different experiments based on three philosophical questions. They are,
@@
  • R1: "How well do the language models perform when only raw evidence sentences along with the claim are available during training and inferencing?"
@@
  • R2: "Does training and inferencing with prompted claim-evidence understanding improve the performance of the language models?
  • R3: "Can training and inferencing with prompted entailed justifications improve the claim veracity prediction?"
Our key contributions in this work are,
@@
  • We conducted three training-based (TBE) and four inference-based (IBEs) experiments along the line of the research questions. We gave three types of inputs, i.e, along with the claim (i) raw evidence sentences, (ii) overall claim-evidence understanding generated by language models, and (iii) entailed justifications generated by the language models.
  • We conducted a detailed evaluation of model explanations. We considered the entailed justifications generated by language models as model explanations, and evaluated them using two strategies: (i) checking lexical overlap and semantic matching, and (ii) doing subjective evaluation by language models. While in the former, we used ROUGE Lin (2004), BLEU Papineni et al. (2002), and BERT-scores Zhang et al. (2019), in the latter, we prompted VLLMs to check how informative, accurate, readable, objective, and logical the explanations are.
  • We conducted an ablation study and a thorough error analysis of our models. In the ablation study, we removed individual supporting and refuting entailed justifications from input samples and trained the veracity prediction models again. This exercise illustrated the importance of individual entailed justifications. In linguistic analysis, on the other hand, we identified (i) the cases where our model succeeds and fails, and (ii) the possible reasons behind it.
Some of the interesting observations we got are,
  • While prompting language models (IBEs) with raw evidence sentences couldn’t outperform the baselines, training some language models with the same (TBE-1) registered an improvement up to 8.20% in macro-F1 for the RAW-FC dataset. Training Language models with prompted claim-evidence understanding (TBE-2) registered an improvement (with a margin up to 16.39%) over baselines for the RAW-FC dataset. Training language models with entailed justifications (TBE-3) outperformed the baselines by a large margin (up to 28.57% and 44.26% for LIAR-RAW and RAW-FC, respectively).
  • Subjective evaluation by language models validated the correlations between veracity prediction and explanation quality in the best-performing models. For instance, Llama-generated explanations received highest informativeness, readability, objectivity, and logicality ratings, which correlates with the highest macro-f1 it got in veracity prediction.
  • In the ablation study, we found that the macro-F1 score dropped when the best performing models were trained without supporting ($\textbf{31.48}\%\downarrow$ in LIAR-RAW and $\textbf{12.50}\%\downarrow$ in RAW-FC) and refuting justifications ($\textbf{9.26}\%\downarrow$ in LIAR-RAW and $\textbf{9.09}\%\downarrow$ in RAW-FC). Further, discarding both had the maximum adverse effect ($\textbf{55.56}\%\downarrow$ in LIAR-RAW and $\textbf{47.73}\%\downarrow$ in RAW-FC).
@@ 2 Dataset details:
In this section, we have reported the details of the datasets considered in our study. We have used the LAIR-RAW and RAW-FC datasets provided by Yang et al. (2022b) in our experiments. While each sample of LAIR-RAW is tagged with one of six labels, samples of RAW-FC is tagged with one of three labels. The list of labels and their distribution in respective datasets are reported in Table 1. The datasets are open-source, i.e., they are Apache 2.0 licensed. A detailed description of the individual datasets along with representative sample is reported in the Appendix B.
@@
Table 1: Datasets Statistics
Figure 1: Illustration of steps we followed in different experiments. Sub-figure (a) presents the case where only raw evidence sentences and claims are given as input. This approach is used in TBE-1 and IBE-1. In TBE-1, we trained the adapters (with VLLMs) using these inputs, whereas in IBE-1, we prompted VLLMs using zero-shot promptings. Sub-figure (b) shows the overall process of TBE-2, IBE-2 and IBE-3. In TBE-2, the training of LLMs and VLLMs with adapters were done with the help of claim-evidence understandings generated by VLLMs. However, in IBE-2 and IBE-3, we used zero-shot and CoT based prompting for final veracity prediction. Sub-figure (c) illustrates the overall experimental process of TBE-3 and IBE-4. Here, first, we generate the entailment labels ("supporting" or "refuting") for individual evidence sentences for a given claim. We prompted the considered VLLMs to do the same. Subsequently, we clubbed the supporting and refuting evidence sentences and prompted VLLMs to generate justifications supporting and refuting the given claim. Lastly, based on the generated justifications, we generated the claim veracity by (i) training the LLMs and VLLMs with adapters as part of TBE-3, and (ii) prompting the VLLMs as a part of IBE-3.
@@ 3 Experiments:
In this section, we have reported the details of the experiments we conducted as part of this study. We conducted two types of experiments- (i) training-based (TBE), and (ii) inference-based (IBE). As their name suggests, in TBE approaches, we finetuned various language models to predict the veracity labels based on customised inputs. Particularly, (i) we fine-tuned large language models (LLMs) like RoBERTa Liu et al. (2019) and XLNet Yang et al. (2019), and (ii) we fine-tuned very large language models (VLLMs) like Mistral Jiang et al. (2023), Llama AI@Meta (2024), Gemma Team et al. (2024), Qwen Yang et al. (2024) and Falcon Almazrouei et al. (2023) models using LoRA Hu et al. (2022) and LoRA+ Hayou et al. (2024) adapters. Note that, unlike LLMs, we can not directly fine-tune VLLMs due to their large parameter size and computational constraints. Similarly, in IBE approaches, we prompted the considered VLLMs to predict veracity labels without explicitly training them for it. We considered the same set of LLMs and VLLMs in all of our experiments reported in this study. The details of individual experiments in each category are reported in the subsequent subsections. We have also reported the current state-of-the-art models as the baselines.
3.1 Training Based Experiments (TBE):
In this section, we have reported the details of three training-based experiments we have considered in our study. Each of them is based on a unique research philosophy. The details of individual approaches are reported in the subsequent subsections.
@@ 3.1.1 TBE-1: Training based on raw-evidences:
In the first experiment, we tried to answer "How well do the language models perform when raw evidence sentences are available during training?". To answer this, we fine-tuned the language models by giving claims and raw evidence sentences as input. For the samples which don’t have associated raw-evidence sentences, we gave claims as the only input. Here, we restricted ourselves from fine-tuning LLMs like RoBERTa Liu et al. (2019) and XLNet Yang et al. (2019), as the input length of many samples exceeded the maximum supported input size of these language models. Past work Cheung and Lam (2023) demonstrated the effectiveness of VLLM finetuning using evidence pieces from web search, whereas in our case, we used the gold evidence sentences given in the dataset. The emergence of adapter-based training for VLLMs allowed us to do this, as they can be trained with large input texts. To the best of our knowledge, we believe we are the first to fine-tune the VLLMs using raw evidence sentences. The approach we followed in TBE-1 is illustrated in sub-figure (a) of Figure 1.
@@ 3.1.2 TBE-2: Training based on overall understanding:
In the second experiment, we tried to answer "Does training with VLLM-generated claim-evidence understanding improve the performance of the language models?. To conduct this experiment, we first prompted the five considered VLLMs to generate their understanding of a given claim and its evidence sentence set. For the samples which don’t have an associated evidence sentence, VLLMs generated their understandings based on the embedded knowledge. Based on the understanding, we fine-tuned the LLMs (such as RoBERTa Liu et al. (2019) and XLNet Yang et al. (2019)) and trained the adapters (LoRA Hu et al. (2022) and LoRA+ Hayou et al. (2024)) with the considered VLLMs to produce the claim veracity. The detailed experimental process is illustrated in sub-figure (b) of Figure 1. Some of the prompt samples are presented in the appendix (Figure 18).
@@ 3.1.3 TBE-3: Training based on entailment understanding:
In the third experiment, we tried to answer "Can training with VLLM-generated entailing justifications improve the claim veracity prediction?". To conduct this experiment, we followed a three-step approach. In the first step, we prompted the considered five VLLMs to classify if the evidence sentences are “supporting" or “refuting" a given claim. Our approach is inspired by the L-Defense model proposed by Wang et al. (2024a), where the authors used an attention mechanism to distinguish supporting and refuting evidence sentences. We, however, used the VLLMs for the same. In the second step, we prompted the language models to generate supporting and refuting justifications based on the classified evidence sentences. For the cases where claims don’t have any supporting or refuting evidence sentences, VLLMs generated justifications based on their embedded knowledge. We separately passed the supporting and refuting evidence sentences associated with a claim in the prompts to generate them. Finally, based on the claims and the generated justifications, we (i) fine-tuned the LLMs (such as RoBERTa Liu et al. (2019) and XLNet Yang et al. (2019)), and (ii) fine-tuned the adapters (LoRA Hu et al. (2022) and LoRA+ Hayou et al. (2024)) with considered VLLMs to generate the veracity labels. The detailed approach is illustrated in sub-figure (c) of Figure 1. Some prompt samples we used at each step are illustrated in Figure 19, Figure 20, and Figure 21 in the appendix.
3.2 Inference Based Experiments (IBE):
We have experimented with four types of inference-based approaches, each based on a unique prompting philosophy. They are (i) zero-shot prompting (IBE-1), where we have prompted five considered VLLMs to predict the veracity label given a claim and its associated evidence sentences, (ii) zero-shot prompting with overall understanding (IBE-2), where, first, we prompted the five considered VLLMs to generate the claim-evidence understanding and then, we prompted them again to with the understanding to predict the veracity labels, (iii) CoT prompting with overall understanding (IBE-3), where we followed similar steps as mentioned in IBE-2, except, additionally, we asked the VLLMs to generate step-by-step reasoning behind their predictions, and (iv) prompting based on entailment (IBE-4), where we prompted the VLLMs to predict veracity labels based on entailed justifications. Due to space constraints, we have reported the details of individual approaches in the appendix. For the cases where claims don’t have associated evidence sentences, we gave the claim as only input (in TBE-1), generated understandings and justifications based on the embedded knowledge of VLLMs.
@@ 3.3 Baselines:
In this section, we have reported the previously proposed best-performing models as the baselines. Particularly, we compared our models with the performances of HiSS Zhang and Gao (2023), FactLLaMa Cheung and Lam (2023), RAFTS Yue et al. (2024), and L-Defence Wang et al. (2024a) models. Out of them, HiSS Zhang and Gao (2023) and FactLLaMa Cheung and Lam (2023) retrieve evidence from external sources, while RAFTS Yue et al. (2024) employs a coarse-to-fine retrieval technique to extract evidence directly from the dataset. In contrast, L-Defense Wang et al. (2024a) used relevant evidence without additional retrieval. The previously reported performance of these models on LIAR-RAW and RAW-FC datasets are presented in Table 2. A detailed description of the individual models is presented in the Appendix C.5.
@@
Table 2: Performance of the considered baseline methods on the LIAR-RAW and RAWFC datasets. Notation: our reproduced results are marked as ‘†’.

### v2

@@ Abstract
Automated fact-checking is a crucial task in this digital age. The NLP community has been trying various strategies to build robust fact-checking systems. However, we have not been very successful yet. One main reason behind this is that fact verification is a complex process. Language models have to parse through multiple pieces of evidence, often contradicting each other, to predict a claim’s veracity. In this paper, we proposed a simple yet effective strategy, where we relied on the entailment ability of language models to improve the fact-checking performance. Apart from that, we did a comparison of different prompting and fine-tuning strategies, as it is currently lacking in the literature. Some of our observations are: (i) training language models with raw evidence sentences (TBE-1) and overall claim-evidence understanding (TBE-2) resulted in an improvement up to 8.20% and 16.39% in macro-F1 for RAW-FC dataset, and (ii) training language models with entailed justifications (TBE-3) outperformed the baselines by a huge margin (up to 28.57% and 44.26% for LIAR-RAW and RAW-FC, respectively). We have shared our code repository to reproduce the results.
@@ 1 Introduction:
Figure 1: A schematic diagram explaining fact-checking based on entailment, which breaks down into three stages. (a) Entailment: Given a claim and its corresponding evidences, a VLLM first distinguishes evidences into supporting/ refuting evidence via entailment. (b) Evidence consolidation: Then the same VLLM consolidates these two groups into concise supporting and refuting justification. (c) Veracity prediction: Using the claim and both justifications, an LLM is trained to predict veracity.
@@
The spread of misinformation on the internet has grown to be a pressing social issue. Its consequences have manifested across social, political and commercial domains Mozur (2018); Fisher et al. (2016); Allcott and Gentzkow (2017); Burki (2019); Aghababaeian et al. (2020). To counter the spread, institutional interventions such as, the International Fact-Checking Network (IFCN) came up, which works with over 170 fact-checking organisations and websites worldwide (as of July 2024  [^1] https://www.poynter.org/ifcn/). However, manually verifying facts and detecting misinformation is a slow and costly process. Also, it is impossible to match the pace at which misinformation evolves and spreads over the internet. To overcome this, the research community have been trying to automate the process. Their interest can be gauged by the fact that more than twelve hundred research articles and fifty datasets were published on this topic Alnabhan and Branco (2024); Guo et al. (2022). A detailed discussion of such works is done in the Appendix A.
@@
Doing automated fact-checking at scale is still an open challenge. Most of the proposed approaches relied on language models Shu et al. (2022); Yang et al. (2022a); Yue et al. (2023); Choi and Ferrara (2024a). Researchers have either (i) relied on their embedded knowledge or (ii) used evidence collected from various sources to predict the veracity. For example, while Pan et al. (2021); Zeng and Gao (2023); Dhuliawala et al. (2024) relied on embedded knowledge by using zero-shot, few-shot and chain-of-thought based approaches, Galitsky (2025); Zhang and Gao (2023) retrieved evidence pieces from the web to verify facts. However, the accuracy of existing systems is not high enough to deploy in the real world. The primary reason behind this is that fact verification is a complex process. Language models have to parse through multiple pieces of evidence, which often contradict each other. The contradictory pieces of evidence confuse the language models, leading to poor performance. To overcome this, we proposed a simple strategy, where we relied on the entailment and the generative ability of language models to improve the fact-checking performance. Particularly, we asked generative very large language models (VLLMs) to produce “supporting” and “refuting” justifications (for the truthfulness of a claim) based on respective pools of evidence sentences, and trained encoder-only language models (LLMs) to produce veracity. This strategy is showcased in Figure 1.
Our approach was inspired by Wang et al. (2024a). They relied on attention mechanisms to extract top-k evidence sentences stating the claim to be ‘true’ and ‘false’. They produced respective justifications by promoting and then trained the language models to produce veracity. We, on the other hand, used VLLMs to classify the sentences as “supporting” or “refuting” and considered all evidence sentences for classification. The simplicity of our approach allows us to easily deploy our system in the real world. Also, in our experiments, we found that our approach improves the overall performance. We also compared our approach with different prompting and fine-tuning strategies through a series of experiments. All of our experiments are done around three research questions. They are,
@@
  • R1: "How well do the language models perform when only claim and raw evidence sentences are available?"
@@
  • R2: "Does prompted claim-evidence understanding improve the performance of the language models?
  • R3: "Can prompted entailed justifications improve the claim veracity prediction?"
Our key contributions and observations in this work can be summarised as follows,
@@
  • We conducted three training-based (TBEs) and four inference-based (IBEs) experiments along the line of the R1, R2 and R3. For RAW-FC, training with raw evidence sentences (TBE-1) and overall claim-evidence understanding (TBE-2) registered an improvement up to 8.20% and 16.39% in macro-F1 over baseline. Training with entailed justifications (TBE-3), on the other hand, outperformed the baselines by a huge margin (up to 28.57% and 44.26% for LIAR-RAW and RAW-FC, respectively).
  • We considered the entailed justifications generated by VLLMs as model explanations, and evaluated them using two strategies: (i) checking lexical overlap and semantic matching, and (ii) doing subjective evaluation by VLLMs itself. Notably, explanations generated by Llama received the highest subjective scores, correlating with the highest registered veracity prediction performance.
  • We conducted (i) an ablation study by removing individual supporting and refuting entailed justifications and (ii) a thorough linguistic analysis of our models. We found that both of the entailed justifications are important, as their removal leads to a decline in $MF$ performance. In linguistic analysis, we found that LLM could give high attention scores to appropriate keywords and factual pieces for samples with ‘true’ and ‘false’ labels. However, for samples with borderline (e.g. half-true) labels, attention scores seem to be scattered.
@@ 2 Dataset details:
In this section, we have reported the details of the datasets considered in our study. We considered the LAIR-RAW and RAW-FC datasets provided by Yang et al. (2022b). We selected them as they are popular and widely used in past works such as HiSS Zhang and Gao (2023), FactLLaMa Cheung and Lam (2023), RAFTS Yue et al. (2024), L-Defence Wang et al. (2024a), and others Wang et al. (2025); Zhang and Gao (2024); Xiong et al. (2025). While each sample of LAIR-RAW is tagged with one of six labels, samples of RAW-FC is tagged with one of three labels. The list of labels and their distribution in respective datasets are reported in Table 1. The datasets are open-source, i.e., they are Apache 2.0 licensed. A detailed description of the individual datasets, along with representative samples, is reported in the Appendix B.
@@
Table 1: Distribution of samples in the LIAR-RAW and RAW-FC datasets.
Figure 2: Illustration of steps we followed in different experiments. Sub-figure (a) presents the case where only raw evidence sentences and claims are given as input (R1). This approach is used in TBE-1 and IBE-1. Sub-figure (b) shows the overall process of TBE-2, IBE-2 and IBE-3 based on R2. Sub-figure (c) illustrates the overall experimental process of TBE-3 and IBE-4 based on R3.
@@ 3 Experiments:
In this section, we have reported the details of the experiments we did in this study. As said earlier, we conducted three training-based (TBEs) and four inference-based (IBEs) experiments. The schematic diagram of individual TBEs and IBEs are illustrated in Figure 2. In TBEs, we finetuned (i) LLMs like RoBERTa Liu et al. (2019) and XLNet Yang et al. (2019), and (ii) VLLMs like Mistral Jiang et al. (2023), Llama AI@Meta (2024), Gemma Team et al. (2024), Qwen Yang et al. (2024) and Falcon Almazrouei et al. (2023) with LoRA Hu et al. (2022) and LoRA+ Hayou et al. (2024) adapters. Note that, unlike LLMs, we can not directly fine-tune VLLMs due to their large parameter size and computational constraints. Similarly, in IBEs, we prompted the considered VLLMs to predict the veracity. The details of individual experiments and the current state-of-the-art baselines are reported in the subsequent subsections.
3.1 Training Based Experiments (TBEs):
@@ 3.1.1 TBE-1: Training based on raw-evidences:
Our first training-based experiment was based on R1, where, we fine-tuned the LLMs by giving claims and raw evidence sentences. If any sample doesn’t have any associated evidence sentence, we gave the claim as only input. We restricted ourselves from using LLMs like RoBERTa Liu et al. (2019) and XLNet Yang et al. (2019), as the length of many input instances exceeded the maximum supported input size. Past work Cheung and Lam (2023) demonstrated the effectiveness of fine-tuning VLLMs using evidence pieces from web search, whereas in our case, we used the gold evidence sentences given in the dataset. The emergence of adapter-based training allowed us to efficiently train the VLLMs with limited computational resources. To the best of our knowledge, we believe we are the first to fine-tune the VLLMs using raw evidence sentences. Our approach is illustrated in sub-figure (a) of Figure 2.
@@ 3.1.2 TBE-2: Training based on overall understanding:
Our second experiment in this line was based on R2. To conduct it, we first prompted the five considered VLLMs to generate their understanding of a given claim and its evidence pieces. For any samples which don’t have an associated evidence sentence, VLLMs generated their understandings based on only the claim. With that, we fine-tuned the LLMs and VLLMs with adapters to produce the claim veracity. The detailed experimental process is illustrated in sub-figure (b) of Figure 2. Some of the prompt samples are presented in the appendix (Figure 18).
@@ 3.1.3 TBE-3: Training based on entailment understanding:
Our final training-based experiment was based on R3. To conduct it, we followed a three-step approach. In the first step, we prompted the considered VLLMs to classify if the associated evidence sentences are “supporting" or “refuting" a given claim. In the second step, we prompted the language models to generate supporting and refuting justifications based on the classified evidence sentences. For the cases where claims don’t have any supporting or refuting evidence sentences, VLLMs generated justifications based on their embedded knowledge. Finally, based on the claims and the generated justifications, we fine-tuned the (i) considered LLMs and (ii) adapters with the VLLMs, to generate the veracity. The detailed approach is illustrated in sub-figure (c) of Figure 2. Some prompt samples we used at each step are illustrated in Figure 19, Figure 20, and Figure 21 in the appendix.
3.2 Inference Based Experiments (IBEs):
We conducted four IBEs based on R1, R2 and R3. They are, (i) zero-shot prompting (IBE-1), where we prompted VLLMs to predict the veracity based on a claim and associated evidence pieces (as per R1), (ii) zero-shot prompting with overall understanding (IBE-2), where, first, we prompted the VLLMs to generate the claim-evidence understanding and then, prompted them again to predict the veracity (as per R2), (iii) CoT prompting with overall understanding (IBE-3), where we followed similar steps as in IBE-2, except, we asked VLLMs to additionally generate the reasoning, and (iv) prompting based on entailment (IBE-4), where we prompted the VLLMs to predict veracity based on entailed justifications (as per R3). Due to space constraints, we have reported the details of individual approaches in the appendix. For the cases where a claim doesn’t have any associated evidence, we gave the claim as the input.
@@ 3.3 Baselines:
In this section, we have reported the previously proposed best-performing models as the baselines. Particularly, we compared our models with the performances of HiSS Zhang and Gao (2023), FactLLaMa Cheung and Lam (2023), RAFTS Yue et al. (2024), and L-Defence Wang et al. (2024a) models. Out of them, HiSS Zhang and Gao (2023) and FactLLaMa Cheung and Lam (2023) retrieved evidence from external sources, while RAFTS Yue et al. (2024) employed a coarse-to-fine retrieval technique to extract evidence directly from the dataset. In contrast, L-Defense Wang et al. (2024a) used relevant evidence without additional retrieval. The previously reported performance of these models on LIAR-RAW and RAW-FC datasets are presented in Table 2. All of the baselines and our experiments have used the same training, validation and test splits provided by Yang et al. (2022b) for justified comparison. A detailed description of the individual models is presented in the Appendix C.5.
@@
Table 2: Performance of the considered baselines for LIAR-RAW and RAWFC datasets. Notation: our reproduced results are marked as ‘†’.